1.An I.V. bag dispenses between 500 ml and 1000 ml of fluid, following a uniform distribution.

a. Draw the distribution below. Include the probability function (the answer to 1b).

b. What is the probability function for the amount of fluids dispensed?

c. What is the probability that the amount of fluids dispensed will be less than 750 ml?

d. What is the probability that the amount of fluid dispensed will be between 500 ml and 1000 ml?

e. What is the probability that the amount of fluid dispensed will be between 600 ml and 800 ml?
Answer #1

Ans:

a)

b)

f(x)=1/(1000-500)=1/500=0.002

c)

P(X<=x)=(x-500)/(1000-500)=(x-500)/500

P(X<=750)=(750-500)/500=0.5

d)

P(500<X<1000)=1

e)

P(600<X<800)=P(X<800)-P(X<600)

=(800-600)/500

=200/500

=0.4